6) Convert the following parametric equation to rectangular equations. x = 2 tant-2 (x = 2 - 2 4 y = 2 + sect . (y = -2
Answer #1

@ Given, x = 2 tanta x+2 = 2 tant =) x+2 = tant sovuaring both sides, (+212 = (tant)? +4+2 (2)X = dan²t y (video on) x² +48+4ye+4-44 -3-48-4 - / fisechf-tun?t=1 4 =) - Hy? +16 - 167 -32-424-4 4y2-22 -164-47 +8=0 This is Rectangular Equation, © t - 2-